# Sharpening Kubernetes Audit Logs with Context Awareness

### v1

@@ 1 Introduction
Kubernetes is a powerful and flexible platform for managing containerized applications. It is designed to be highly scalable, allowing users to easily add or remove nodes and workloads as needed. However, scaling up a K8s cluster dramatically increases the complexity of the system and the amount of data it generates. Despite this added complexity, users’ requirements and expectations for performance and reliability do not change. As a result, the need for efficient monitoring and diagnosis of issues in K8s clusters has become increasingly important [1].
@@
Nothing makes a point clearer than a concrete example: in a controlled experiment using a three-node Kubernetes cluster (details follow in later sections), we observed a considerable level of complexity in the system’s behavior. We ran 238 distinct K8s actions, executed a total of 1155 times. Among these, one of the most common, and practically unavoidable, was the creation of a Deployment: a K8s object responsible for declaratively managing replicated pods, including automated rollouts and rollbacks. Every time this action was triggered (what we refer to as the triggering event), K8s audit produced a bulky log line, like the one shown in Figure 1, containing 4700 characters. In addition to that, it produced, on average, another 16 similarly large lines, describing secondary actions related to it. These additional events are often neither directly linked together nor uniquely associated with the Deployment action alone; indeed, many of them can be triggered by different actions across the system. In our contained experimental setting, Deployment actions occurred 40 times and resulted in 636 log lines. If that does not already suggest a tangled puzzle, consider this: deleting a Namespace generates even more complexity, as each resource within the Namespace must first be assessed and then individually removed. In just three Namespace deletions, our system produced 408 log lines, all unrelated to the original triggering event.
@@ 2.1 Kubernetes
Kubernetes (K8s[5] is an open-source container orchestration platform that automates the deployment, scaling, and management of containerized applications. It was originally developed by Google, and it is now maintained by the Cloud Native Computing Foundation (CNCF). K8s is designed to be extensible and scalable, allowing users to deploy and manage containerized applications across a cluster of machines.
@@ 3.3.1 Malfunctioning components
As an example, following a successful upgrade of a Kubernetes cluster, the control plane failed to establish secure communication with its nodes. Although the nodes themselves remained healthy—continuing to send heartbeats, the leader election process repeatedly timed out, and the control plane could not schedule or deploy any new workloads. Inspection of the extracted log entries revealed that every attempt by the kube-controller-manager to fetch its own Lease object under the coordination.k8s.io API was being met with HTTP 401 "Unauthorized" responses. This kind of problem is not easy to identify, and even during a cluster’s boot-up phase, the audit logging module will generate thousands of lines. This breakdown in TLS certificate validation effectively severed the control plane–node channel and halted cluster operations.
@@ 3.3.3 Evading detection
Assume the attacker has successfully leaked a user’s credentials in the cluster. Even if a security expert suspects that a user has been compromised, looking specifically at the audit log is not straightforward since it only shows individual API calls without enough context to tell routine admin work from malicious activity. On top of that, attackers often stick to valid API calls or built-in escalation paths that raise no errors. Hence, the administrator needs to manually piece together behavioral patterns, network indicators, and threat intelligence just to hint at the real compromise path.
@@ 5.3 Log parser
To begin with, audit logging must be properly configured in the K8s cluster [2]. Kubernetes auditing natively supports several criteria for filtering events at the source, such as the resource type or the user that performed the action. However, in our implementation, no filtering is performed. This allows K8NTEXT to ingest all events generated by the cluster, and thus, to have a comprehensive view of the cluster’s activity and to truly capture every dependency and interaction between events. Filtering at the source would reduce the amount of data ingested, but it would also provide K8NTEXT with an incomplete view of the cluster’s activity, potentially leading to missed dependencies and interactions between events.
@@ 5.5 Labels
Then, we augment this information by introducing two additional fields, objectRef.namespace and objectRef.name, that are present in the logs only if the object exists in a namespace and the action was performed on a single object, respectively. 666 K8s differentiates between namespaced and non-namespaced objects, which can affect how events are processed and interpreted. The latter are cluster-wide objects, while the former are confined to a specific namespace. Some resources may exist in both namespaced and non-namespaced versions. Second, some actions can be performed on a single object, while others can be performed on multiple objects.
@@ 5.8.1 Input and output shapes
The second and third ones, on the other hand, determine the output shape of the model. As described in Section 5.5, the model is trained to predict a label that is a tuple of integers. This tuple is composed of $|L|$ elements, and each element is an integer that, once encoded, can take a value between 00 and $m-1$. Our model outputs a tensor of shape $(W,|L|,m)$: in other words, for each log line in the batch ($W$), we can reconstruct the original tuple $L$ by applying a softmax activation function to each element $L_{i}$ of the output tensor.
@@ 5.9 Clustering
At the end of the labeling process, each log line has been assigned a label that uniquely identifies the triggering event it is part of. However, recall that we explicitly embedded only a subset of the fields in the label, namely, the generic fields. This means that while triggering events and their consequences are categorized, the specific contexts in which they occur are not yet fully reconstructed. As an example, imagine the creation of two namespaces. Their label will be shared, but the contexts in which they were created will be different. At the moment, we have no way to distinguish between them.
To do so, we need to group log lines that are related to the same triggering event together. This is a non-trivial task, as the log lines are not ordered in any way, and the only information we have is the label of the triggering event. We can pursue different clustering approaches.
5.9.1 Matching objectRef
We can group log lines together that have the same objectRef field. This is a simple approach that is guaranteed to work if the object is namespaced and single.
5.9.2 Matching involvedObject, claimRef, ownerReferences
Some advanced K8s objects, such as Deployments, StatefulSets, or Events generate a cascade of objects and events when they are created, updated, or deleted. For example, creating a Deployment called test will first create a ReplicaSet, called test-xxxxx, which will then create one or more Pods, called test-xxxxx-yyyyy. The linking information is usually found in the request or response bodies in the form of the ownerReferences field. A similar role is played by the claimRef field in PersistentVolumeClaims and PersistentVolumes, and the involvedObject field in Events. By matching these fields, we can group log lines together that are related to the same object.
5.9.3 Uniform distribution
If no other field is present that can assist in the clustering process, but a certain set of log lines is present in the same quantity or multiple of the triggering event(s), we can divide them among the contexts based on their occurrence. For example, the act of describing a Deployment is internally translated into a series of get and list operations on all pods and events related to it. Assume for simplicity that each kubectl describe executes one get and one list. These actions are usually executed by the user itself (and in this case, linking them is trivial), but sometimes by the control plane proxies it or fetches other contextual information to complete the user requests. If two users are both describing the same Deployment, we will have two log lines for the get action and two log lines for the list action. If we cannot match them by user, we can assign the first log line to the first context, the second to the second context, and so on. This is a simple approach that works well in most cases that have a uniform distribution of log lines and no clear linking information.
@@ 5.10 Querying
When enabled, users can query the contexts that have been reconstructed using a simple query language. Using classic boolean operators such as ,,<,,$\land$,$\lor$,<$\geq$,$\lnot$ , ∨ , < , ≥ , ¬, and the exists and regexp operators, users can express complex queries that are then parsed by the engine. For example, the query verb == "create" and objectRef.namespace == "default" will return all the contexts that are related to the creation of an object in the default namespace.
@@ 6.2 Window size
The model’s training time scales linearly with the window size, but the size on disk does not. This is because the model’s learnable parameters are not tied to the window size, and the model’s size on disk remains constant, regardless of the window size: $\approx$ 718,000 parameters or 2.7 MB. Such a dimension is acceptable for a deep learning model that needs to perform real time inference on data on a standard cluster; however, the performance may suffer on more resource-constrained scenarios.
The model’s inference time also scales linearly with the window size, as the model needs to process more log lines at once. We measured the inference time at the same $W$ values as the training time and using the same test dataset used for training. The results are shown in Table 6. The total inference time is the sum of the model inference time and the time needed to decode the labels divided by the total number of log lines in the batch, measured in milliseconds. The standard deviation is not reported as it is negligible in this case.
@@ 6.5.1 Dataset splits
To assess this claim, we employed a k-fold cross-validation approach to verify how the accuracy varies across the dataset, with increasing values of $k$. To be even more demanding in our evaluation, we used majority accuracy as mentioned before. Majority accuracy takes into account how many log lines were correctly predicted, i.e., correct linestotal $\frac{\text{correct lines}}{\text{total lines}}$, after batches are flattened and labels are assigned back to lines. This allowed us to evaluate the model’s performance on different splits of the dataset and truly obtain an accurate estimate of the model’s performance on unseen data.
@@ 7 Capability and usability evaluation
This section evaluates K8NTEXT’s performance in terms of clustering rate, querying time, and storage space. These metrics are important for assessing the model’s usability and efficiency in real-world scenarios. All the experiments were run under the same conditions as the previous ones, using the same virtual machine and dataset.
7.1 Clustering rate
@@
The end goal of K8NTEXT is to cut through the noise of the logs, providing users with a manageable set of contextualized events that can be used to understand the system’s behavior. To achieve this, K8NTEXT clusters log lines together based on the labels assigned by the model and then further subdivided as described in Section 5.9. The clustering rate is a measure of how well the model can do this, thus reducing the overall number of events that need to be inspected.
Figure 10 summarizes the results of this experiment. 45% of the labels cluster 1 to 4 log lines together, representing quick and simple actions, such as listing a resource. On the other side of the spectrum, a handful of labels cluster 100 or more lines together: for example, the update of a Deployment. This action is a very complex one, which involves updating the underlying ReplicaSets and its Pods. Another mammoth-sized action is the deletion of a Namespace: this action is very complex and involves deleting all the resources contained in the Namespace, which is usually hundreds of log lines.
Overall, K8NTEXT shows a very good clustering rate: more than 50% of the labels cluster 5 or more lines together, 16 labels cluster 50 or more lines together, and 4 labels cluster 100 or more lines together. This means that K8NTEXT is able to reduce the number of log lines that need to be inspected by a significant amount.

### v2

@@ 1 Introduction
K8s is a powerful and flexible platform for managing containerized applications. It is designed to be highly scalable, allowing users to easily add or remove nodes and workloads as needed. However, scaling up a K8s cluster dramatically increases the complexity of the system and the amount of data it generates. Despite this added complexity, users’ requirements and expectations for performance and reliability do not change. As a result, the need for efficient monitoring and diagnosis of issues in K8s clusters has become increasingly important [1].
@@
Nothing makes a point clearer than a concrete example: in a controlled experiment using a three-node K8s cluster (details follow in later sections), we observed a considerable level of complexity in the system’s behavior. We ran 238 distinct K8s actions, executed a total of 1155 times. Among these, one of the most common, and practically unavoidable, was the creation of a Deployment: a K8s object responsible for declaratively managing replicated pods, including automated rollouts and rollbacks. Every time this action was triggered (what we refer to as the triggering event), K8s audit produced a bulky log line, like the one shown in Figure 1, containing 4700 characters. In addition to that, it produced, on average, another 16 similarly large lines, describing secondary actions related to it. These additional events are often neither directly linked together nor uniquely associated with the Deployment action alone; indeed, many of them can be triggered by different actions across the system. In our contained experimental setting, Deployment actions occurred 40 times and resulted in 636 log lines. If that does not already suggest a tangled puzzle, consider this: deleting a Namespace generates even more complexity, as each resource within the Namespace must first be assessed and then individually removed. In just three Namespace deletions, our system produced 408 log lines, all unrelated to the original triggering event.
@@ 2.1 Kubernetes
Kubernetes (K8s) [5] is an open-source container orchestration platform that automates the deployment, scaling, and management of containerized applications. It was originally developed by Google, and it is now maintained by the Cloud Native Computing Foundation (CNCF). K8s is designed to be extensible and scalable, allowing users to deploy and manage containerized applications across a cluster of machines.
@@ 3.3.1 Malfunctioning components
As an example, following a successful upgrade of a K8s cluster, the control plane failed to establish secure communication with its nodes. Although the nodes themselves remained healthy—continuing to send heartbeats, the leader election process repeatedly timed out, and the control plane could not schedule or deploy any new workloads. Inspection of the extracted log entries revealed that every attempt by the kube‐controller‐manager to fetch its own Lease object under the coordination.k8s.io API was being met with HTTP 401 "Unauthorized" responses. This kind of problem is not easy to identify, and even during a cluster’s boot-up phase, the audit logging module will generate thousands of lines. This breakdown in TLS certificate validation effectively severed the control plane–node channel and halted cluster operations.
@@ 3.3.3 Evading detection
Assume the attacker has successfully leaked a user’s credentials in the cluster. Even if a security expert suspects that a user has been compromised, looking specifically at the audit log is not straightforward since it only shows individual API calls without enough context to tell routine admin work from malicious activity. On top of that, attackers often stick to valid API calls or built‑in escalation paths that raise no errors. Hence, the administrator needs to manually piece together behavioral patterns, network indicators, and threat intelligence just to hint at the real compromise path.
@@ 5.3 Log parser
To begin with, audit logging must be properly configured in the K8s cluster [2]. K8s auditing natively supports several criteria for filtering events at the source, such as the resource type or the user that performed the action. However, in our implementation, no filtering is performed. This allows K8NTEXT to ingest all events generated by the cluster, and thus, to have a comprehensive view of the cluster’s activity and to truly capture every dependency and interaction between events. Filtering at the source would reduce the amount of data ingested, but it would also provide K8NTEXT with an incomplete view of the cluster’s activity, potentially leading to missed dependencies and interactions between events.
@@ 5.5 Labels
Then, we augment this information by introducing two additional fields, objectRef.namespace and objectRef.name, that are present in the logs only if the object exists in a namespace and the action was performed on a single object, respectively.  [^6] K8s differentiates between namespaced and non-namespaced objects, which can affect how events are processed and interpreted. The latter are cluster-wide objects, while the former are confined to a specific namespace. Some resources may exist in both namespaced and non-namespaced versions. Second, some actions can be performed on a single object, while others can be performed on multiple objects.
@@ 5.8.1 Input and output shapes
The second and third ones, on the other hand, determine the output shape of the model. As described in Section 5.5, the model is trained to predict a label that is a tuple of integers. This tuple is composed of $|L|$ elements, and each element is an integer that, once encoded, can take a value between 0 and m1m-1. Our model outputs a tensor of shape $(W,|L|,m)$: in other words, for each log line in the batch ($W$), we can reconstruct the original tuple $L$ by applying a softmax activation function to each element $L_{i}$ of the output tensor.
@@ 5.9 Clustering
At the end of the labeling process, each log line has been assigned a label that uniquely identifies the type of triggering event it is part of. However, we explicitly embedded only a subset of the fields in the label, namely, the generic fields. This means that while triggering events and their consequences are categorized by type, the specific contexts in which they occur are not yet fully reconstructed. As an example, imagine the creation of two namespaces. Their label will be shared, but the contexts in which they were created will be different. At the moment, we have no way to distinguish between the individual contexts for each namespace creation.
To address this, we need to group log lines that are related to the same specific triggering event instance together. This is a challenging task: log lines may arrive out of order, timestamps may be imprecise due to the distributed nature of K8s, and the linking information between events may be scattered across different fields in the request and response bodies. Moreover, in production environments with high operational frequency and many concurrent users, the clustering algorithm must be robust enough to handle large volumes of log lines without incorrectly assigning events to the wrong contexts.
5.9.1 Windowing strategy
To ensure scalability and reliability, the clustering algorithm operates on time-bounded windows of log lines rather than processing the entire log stream at once. This approach mirrors real-world scenarios where audit logs are sent in batches to webhook endpoints. K8s uses a default cutoff of 10,000 lines or 30 seconds (whichever comes first) when sending audit events to webhooks.
Our implementation introduces two configurable cutoff based on the aforementioned K8s parameters: a time threshold and a set maximum lines per window. For each label group identified by the model, we partition the log lines into windows based on these constraints. A new window begins when either the time elapsed since the window’s first line exceeds the time threshold or the window reaches the maximum number of lines. This dual-threshold approach prevents both an unbounded growth of clusters and the accumulation of stale data in low-traffic clusters.
Within each window, the clustering algorithm analyzes all lines before making assignment decisions, ensuring that contextual information is properly considered rather than relying on simple sequential assignment strategies. While this windowing approach may introduce errors at the boundaries (e.g., if related events fall into different windows), it provides a practical balance between performance and accuracy, especially in high-throughput environments.
5.9.2 Cluster assignment
The core of the clustering algorithm is a scoring mechanism that evaluates how well each log line matches existing clusters within the current window. When processing a window, the algorithm first identifies any triggering actions present, i.e., log lines whose label matches their intrinsic characteristics, indicating they directly caused the action rather than being consequences of another event. Each triggering action becomes the seed of a new cluster.
For each remaining log line in the window, the algorithm computes a compatibility score (a simple integer, $s$) with every existing cluster. The score is based on multiple weighted criteria that capture different types of relationships between K8s objects and events. The line is then assigned to the cluster with the highest score, provided it exceeds a minimum threshold ($s=60$). If no cluster scores above the threshold, a new cluster is created.
The scoring hierarchy can be summarized as follows, in decreasing order of importance:
  1. 1.
    ObjectRef match: Direct reference to the same resource (namespace, name, type) is the strongest indicator of relatedness.
  2. 2.
    UID relationships: If a log line’s object UID matches an ownerReferences UID in the cluster’s metadata, or vice versa, this indicates a parent-child relationship. For example, a Deployment creates ReplicaSets that create Pods, and the UIDs link these hierarchically.
  3. 3.
    InvolvedObject match: Events (e.g., Kubernetes Event resources) reference other objects via involvedObject. Matching this field or matching an involved object to a resource in the cluster indicates correlation.
  4. 4.
    ClaimRef match: PersistentVolumes and PersistentVolumeClaims are linked via claimRef, and matching this field groups related storage events.
Username and verb are also matched, but they are considered secondary and insufficient to warrant clustering on their own. Additionally, the algorithm includes targeted heuristics for known K8s behaviors. For instance:
  • 1.
    When a user creates a resource in a namespace, the API Server often performs a list operation on limitranges in that namespace. The algorithm detects this pattern and assigns these system actions to the user’s context with a high score (90 points).
  • 2.
    When a namespace is deleted, the system performs deletecollection and get operations on all resources within that namespace. The algorithm identifies these cascading actions and groups them with the namespace deletion context (95 points).
5.9.3 Control plane and special cases
Some clusters may benefit from merging all control plane actions (actions performed by system components like the API server and the Kube Controller Manager, etc.) into a single dedicated context. This is controlled by a configurable parameter. When enabled, all control plane actions are assigned a fixed UUID, allowing auditors to separate user-initiated actions from background system operations.
The clustering algorithm also handles edge cases gracefully. For instance, single-line windows are trivially assigned to their own cluster. Windows with no triggering actions distribute lines based on the best scoring matches, creating new clusters as needed. The explicit decision hierarchy ensures that high-confidence matches (e.g., ObjectRef or UID relationships) take precedence over lower-confidence heuristics.
@@ 5.10 Querying
When enabled, users can query the contexts that have been reconstructed using a simple query language. Using classic boolean operators such as $\land,\lor,<,\geq,\lnot$, and the exists and regexp operators, users can express complex queries that are then parsed by the engine. For example, the query verb == "create" and objectRef.namespace == "default" will return all the contexts that are related to the creation of an object in the default namespace.
@@ 6.2 Window size
The model’s training time scales linearly with the window size, but the size on disk does not. This is because the model’s learnable parameters are not tied to the window size, and the model’s size on disk remains constant, regardless of the window size: $\approx$ 718,000 parameters or 2.7 MB. Such a dimension is acceptable for a deep learning model that needs to perform real-time inference on data on a standard cluster; however, the performance may suffer in more resource-constrained scenarios.
The model’s inference time also scales linearly with the window size, as the model needs to process more log lines at once. We define the inference time as the total time needed to process a batch of log lines, including all the data encoding and decoding, as well as the majority voting process. This time is then averaged over all the log lines in the batch, and includes the one-time overhead of loading the model in memory. We measured the inference time at the same $W$ values as the training time and using the same test dataset used for training, fed all at once. The results are shown in Table 6. Standard deviation is not reported as it is negligible in this case.
@@
Finally, to ensure that the batch size does not affect these results, we repeated the experiment with different batch sizes, ranging from 500 to 5000 log lines. The results showed that for reasonable batch sizes (i.e., 2000 or more log lines, per the default K8s webhook parameters), the overhead of loading the model in memory is amortized over the entire batch, leading to a negligible difference in inference time per line (less than 0.05 ms with $B=2000$ compared to the previous result).
To conclude, K8NTEXT shows good performance both in terms of training time and inference time, even with moderate window sizes. As re-training is needed when $W$ is changed – a limitation of the current approach – users should carefully evaluate the best window size for their use case before deploying K8NTEXT in production.
@@ 6.5.1 Dataset splits
To assess this claim, we employed a k-fold cross-validation approach to verify how the accuracy varies across the dataset, with increasing values of $k$. To be even more demanding in our evaluation, we used majority accuracy as mentioned before. Majority accuracy takes into account how many log lines were correctly predicted, i.e., $\frac{\text{correct lines}}{\text{total lines}}$, after batches are flattened and labels are assigned back to lines. This allowed us to evaluate the model’s performance on different splits of the dataset and truly obtain an accurate estimate of the model’s performance on unseen data.
@@ 7 Capability and usability evaluation
This section evaluates K8NTEXT’s performance in terms of clustering effectiveness, querying time, and storage space. These metrics are important for assessing the model’s usability and efficiency in real-world scenarios. All the experiments were run under the same conditions as the previous ones, using the same virtual machine and dataset.
7.1 Clustering
@@
The end goal of K8NTEXT is to cut through the noise of the logs, providing users with a manageable set of contextualized events that can be used to understand the system’s behavior. To achieve this, K8NTEXT clusters log lines together based on the labels assigned by the model and then further subdivides them as described in Section 5.9.
To establish a ground truth, we ran our clustering algorithm on the original labels in the dataset, which are, to the best of our knowledge, correct. The clustering algorithm yielded 9843 clusters, with an average of 1.87 log lines per cluster. Since each label produces distinct clusters, we manually inspected a subset of significant labels to ensure that the clustering algorithm was working as intended. Once we were satisfied with the results, we then evaluated the clustering performance of the algorithm both on the original labels and those predicted by the model.
The clustering rate is a measure of how efficiently the model can join lines together, by reducing the overall number of events that need to be inspected. Figure 10 summarizes the results of this experiment. 92 out of 238 labels cluster 2 to 4 log lines together, representing quick and simple actions, such as listing a resource. On the other side of the spectrum, a handful of labels cluster 20 or more lines together: for example, the update of a Deployment. This action is a very complex one, which involves updating the underlying ReplicaSets and their Pods. Another mammoth-sized action is the deletion of a Namespace: this action is very complex and consists in deleting all the resources contained in the Namespace, which is usually hundreds of log lines. Finally, 122 labels cluster on average only a single log line, representing simple actions that do not require any context to be understood, such as the creation of a RoleBinding, or a get action on a Node.
Figure 11: Clustering distribution comparison of the ground truth (blue, light) and the predicted labels (red, dark). Clusters are sorted by their size, and only the last clusters are shown for brevity. The x-axis is shown starting from $x=8500$ for readability.
By itself, the clustering rate is not a sufficient metric to evaluate K8NTEXT’s performance. Thus, we also evaluated the quality of the clusters produced by the model. We ran several instances of the model using a k-fold-like approach as discussed in Section 6.5, merged the predictions, and ran the clustering algorithm on top of them. The clustering algorithm yielded 9756 clusters, 87 less than the ground truth. Of these, 9744 (99.8%) percent were perfectly matched with the ground truth, while the remaining were mixed matches, merges, or splits. Figure 11 shows the distribution of the two clustering results, sorted by their size.
@@ 7.3 Storage space
7.4 Applicability to other scenarios
Figure 12: Timeline of actions performed by different users in Scenario 1. Each color represents a different user, and each block represents an action performed by that user. ns is short for Namespace.
Figure 13: Timeline of actions performed by different users in Scenario 2. Each color represents a different user, and each block represents an action performed by that user. ns is short for Namespace and np is short for NetworkPolicy.
Finally, we evaluated how well K8NTEXT performs in some alternative scenarios that differ from our training dataset. Using a deterministic scenario simulator, we generated two scenarios that mimic real-world use cases. The scenarios were designed to test K8NTEXT’s ability to handle different types of workloads and user behaviors.
Scenario 1 (shown in Figure 12) involves three users concurrently creating, deleting, and scaling Deployments. This scenario has been specifically chosen as a worst-case scenario for K8NTEXT because the actions performed by the users are very similar and occur in close proximity, making it difficult to distinguish between them. On the other hand, Scenario 2 (shown in Figure 13) involves the application of NetworkPolicies and the creation of Pods to verify network connectivity.
Since these datasets are unlabeled, we evaluated K8NTEXT’s performance by running the model, clustering, and extracting the key actions performed by each user and comparing them with the list of user actions shown in Figures 12 and 13. In Scenario 1, of the 11 distinct actions that are performed, 6 are correctly detected (all the Deployment creations and deletions) and the context is reconstructed, 2 are correctly clusterized but not tagged as triggering actions, and 3 are mislabeled and end up in different clusters. In Scenario 2, all the actions are correctly detected and clusterized, with only one mislabeled action. Figures 12 and 13 show a tick ($\checkmark$) next to the correctly identified actions, a cross ($\times$) next to the mislabeled actions, and an approximate sign ($\approx$) next to lines clustered without being tagged as triggering actions.
Overall, K8NTEXT shows a good performance in these scenarios, correctly identifying and clustering most of the actions performed by the users. While there is room for improvement in terms of labeling accuracy, the clustering algorithm is able to reconstruct the context of the actions performed by the users, providing a clear and concise view of the system’s behavior.